Shift report 20010817 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Aug 17 2001 - 00:05:46 EDT

  • Next message: Apache: "Shift report 20010817 00:00-08:00"

    ------------------------------------------------------------
    Date: Fri Aug 17 00:05:45 2001, Shift: 16:00-24:00
    Supervisor: sjs, with: ykl
    Log book#7 pages 102 - 106
    
    Shift Summary:
    ------------------------------------------------------------
    Good beam after 19:30.
    
    Blown fuse in si5_4 FERA replaced.
    FFS and BFS moved to 4 deg--No serious complaints by TPC's.
    (RD expresses conditional satisfaction in log book...)
    Low voltage for T5 had to be recycled.
    
    Run 4245 started at 20:47 and still running. Currently has 
    181K trigger 1 events.
    
    Scan of Supermon plots (of which there are too many!)
    shows C1 ADC1 dead, H1 TDC Bot33 missing 
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Fri Aug 17 2001 - 00:06:16 EDT